Fashionisto Inc. is an upscale clothing store in New York City and London. Each store has two main departments, Men’s Apparel and Women’s Apparel. Marie Phelps, Fashionisto’s CFO, wants to use strategic performance measurement to better understand the company’s financial results. She has decided to use the profit center method to measure performance and has gathered the following information about the two stores and the two departments of the New York City store:

Total net sales $ 4,200,000
Fixed costs
Partly traceable and controllable 285,000
Partly traceable but noncontrollable 255,000
Nontraceable costs 105,000
Total net sales (percent)
London store 40 %
New York store 60
New York—Men’s Apparel 40
New York—Women’s Apparel 60
Cost of goods sold—variable (percent of sales)
London store 52 %
New York—Men’s Apparel 45
New York—Women’s Apparel 60
Variable operating costs (percent of sales)
London store 36 %
New York—Men’s Apparel 22
New York—Women’s Apparel 32
Fixed controllable costs—partly traceable (percent of total)
London store 40 %
New York total 40
Men’s Apparel 45
Women’s Apparel 40
Could not be traced to Men’s or Women’s Apparel 15
Could not be traced to London or New York 20
Fixed noncontrollable costs—partly traceable (percent of total)
London store 50 %
New York total 40
Men’s Apparel 35
Women’s Apparel 15
Could not be traced to Men’s or Women’s Apparel 50
Could not be traced to London or New York 10

Required:

1. Using this information, prepare a contribution income statement for Fashionisto.
Answer #1

FASHIONISTO

CONTRIBUTION MARGIN INCOME STATEMENT

Total Other London New York New York
Total Men's Apparel Women's Apparel Other
Sales 4200000 4200000*40%=16800000 4200000*60%=2520000 2520000*40%=1008000 2520000*60%=1512000
Variable expense
Variable cost of sales 2234400 16800000*52%=873600 1360800 1008000*45%=453600 1512000*60%=907200
Variable operating cost 1310400 16800000*36%=604800 705600 1008000*22%=221760 1512000*32%=483840
Total variable expense 3544800 1478400 2066400 675360 1391040
contribution margin 655200 201600 453600 332640 120960
Traceable Fixed cost
Controllable and traceable 285000 285000*20%=57000 285000*40%=114000 285000*40%=114000 114000*45%=51300 114000*40%=45600 114000*15%=17100
Non controllable and Traceable 255000 255000*10%=25500 255000*50%=127500 255000*40%=102000 102000*35%= 35700 102000*15%=15300 102000*50%=51000
Total Traceable fixed cost 540000 82500 241500 216000 87000 60900 68100
segment margin 115200 -82500 -39900 237600 245640 60060 -68100
Less:Non Traceable fixed cost 105000
Net operating income 10200